Ok here's my question..... The length of a rectangle is 3yards more than twice it's width. The perimeter of the rectangle is 36 yards. What are the width and length of the rectangle? It seemed so easy in class but now I'm having troubles how can I solve problems like these in the future? Please Excuse the E-mail address =)
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Let represent the width. Then represents the length. The perimeter is two times the length plus two times the width, so:



Solve for . Then calculate .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

The length of a rectangle is 3 yards more than twice it's width. The perimeter of the rectangle is 36 yards. What are the width and length of the rectangle?
-----------------------
Let the width be "x".
Then the length = "2x+3".
-------------------------
Perimeter = 2(length + width)
36 = 2(2x+3 + x)
18 = 3x+3
3x = 15
x = 5 yds (width)
---
2x+3 = 2*5+3 = 13 yds (length)
=================================
Cheers,
Stan H.